### v1

# Randomized Masked Finetuning: An Efficient Way to Mitigate Memorization of PIIs in LLMs

Abstract
The current literature on memorization in Natural Language Models, especially Large Language Models (LLMs), poses severe security and privacy risks, as models tend to memorize personally identifying information (PIIs) from training data. We introduce Randomized Masked Fine-Tuning (RMFT), a novel privacy-preserving fine-tuning technique that reduces PII memorization while minimizing performance impact. Using the Enron Email Dataset, we demonstrate that RMFT achieves an 80.81% reduction in Total Extraction Rate and 80.17% reduction in Seen Extraction Rate compared to baseline fine-tuning, outperforming deduplication methods while maintaining only a 5.73% increase in perplexity. We present MaxTER, a Pareto-optimal evaluation framework for assessing privacy-utility tradeoffs, and show the performance of RMFT vs Deduplication by Area Under The Response Curve (AURC) metric.
Machine Learning, Privacy, Large Language Models, Memorization, PII Protection
1 Introduction
The current literature on memorization in Natural Language Models, especially Large Language Models (LLMs), poses a severe security and privacy risk  (carlini2021extractingtrainingdatalarge; yang2024memorization), as models tend to memorize pieces of personally identifying information (PIIs) that occur in pre-training or fine-tuning datasets and can later regurgitate the exact PIIs when prompted. This effect arises from models’ ability to identify token sequences in PIIs that occur frequently, which leads to memorization of the exact PII as compared to generalization and generation when prompted. This was explored by Kandpal and team (kandpal-etal-2024-user) in their study on deduplication training and how it mitigates privacy risks. They and others  (10.1162/TACL.a.49) have observed that token sequences that were duplicated multiple times in the dataset also tend to be memorized more often than token sequences that appeared a smaller number of times or only once.
Our preliminary examination on a GPT2-XL tokenizer revealed that the tokenizer considered a PII as a composition of several tokens. This means that PIIs that occurred multiple times tend to be memorized as a token sequence by the LLM. Therefore, we move forward with the assumption that the N-Gram appearances of PIIs directly impact the ability of an LLM to memorize the PII.
Keeping this in mind, we devised a novel training technique that reduces the number of times a PII appears in the dataset while keeping the original PII intact for future downstream tasks. This technique is called Randomized Masked Fine-Tuning (RMFT). As the name suggests, we mask the PII appearances with a randomized token sequence generation that closely mimics the PII being masked. Using this method, we were able to reduce the number of appearances of certain PIIs from thousands down to a single appearance.
We fine-tune the model with this modified dataset and then test the updated model on the same prompts. We show that RMFT is highly efficient in reducing memorization in LLMs compared to baseline fine-tuning with no dataset modifications, with a minor impact on model performance as measured using perplexity on the WikiText2 benchmark dataset.
We then compare our technique with the deduplication technique proposed by Katherine Lee and team. (lee2022deduplicating) and create a comprehensive suite of evaluations on PII extraction, model performance, and training efficiency. Using our results, we show how RMFT outperforms deduplicated models when it comes to PII extraction on the same prompts.
1.1 Contributions
This study introduces:
  • Randomized Masked Fine-tuning (RMFT): A novel, privacy-preserving fine-tuning technique designed to minimize performance impact while significantly reducing PII memorization.
  • Comprehensive evaluation metrics: Including Total Extraction Rate (TER) and Seen Extraction Rate (SER) for measuring memorization, and Mean Delta Perplexity (MDP) for performance assessment.
  • MaxTER: A new optimization function using Pareto frontier analysis to evaluate the memorization-versus-performance tradeoff.
  • Empirical validation: Demonstrating RMFT’s superiority over deduplication on the Enron Email Dataset with GPT2-XL.
2 Related Work
2.1 PII Memorization and Leakage
Prior work on Personally Identifiable Information (PII) memorization and leakage has revealed a critical, inherent privacy vulnerability in large language models (LLMs). Multiple earlier studies successfully retrieved training data, including PII, from publicly available GPT-2 models (carlini2021extractingtrainingdatalarge; nasr2023scalableextractiontrainingdata; schwarzschild2024rethinkingllmmemorizationlens). Huang and team (huang-etal-2022-large) formally coined the term memorization in 2022 during their work on extracting personal information from pre-trained LLMs. Borkar’s study (borkar2023can) worked extensively on data leakage related to law-based datasets and accredited unlearning as a possible solution to memorization. Extending this concept, Borkar ad team  (borkar2025privacyrippleeffectsadding) introduced the detailed concepts of assisted memorization, retained memorization, forgotten memorization, and immediate memorization. Furthermore, multiple studies have focused on the memorization of token sequences, phrases, and passages and the capacity of LLMs to reproduce them exactly as available in the training data (satvaty2025undesirablememorizationlargelanguage; hartmann2023sokmemorizationgeneralpurposelarge; wang2025generalizationvsmemorizationtracing).
2.2 Mitigation Techniques
There are multiple mitigation techniques that aim at reducing the memorization of PIIs and token sequences in LLMs while maintaining minimal impact on model performance. Unlearning (borkar2025privacyrippleeffectsadding; chen2023unlearnwantforgetefficient; si2023knowledgeunlearningllmstasks; tamirisa2024toward) and deduplication (kandpal-etal-2024-user; lee2022deduplicating) are often cited as the most efficient mitigation techniques currently. However, a study by To and Le  (to2025harrypotterhereprobing) indicated that unlearning can be inefficient and introduced a framework called LURK, which evaluates unlearned models for retained memorization of unlearned data. Kandpal  (pmlr-v162-kandpal22a) accredited differential privacy as an efficient mitigation technique for membership inference. In a comparative study, Sakarvadia and team  (sakarvadia2025mitigatingmemorizationlanguagemodels) compared 11 different techniques used to combat memorization and introduced balanced subnet, a model modification-based method. Separately, Suri and team  (suri2025mitigatingmemorizationllmsusing) proposed activation steering on privacy neurons as an efficient mitigation method.
2.3 Privacy vs Utility in LLMs
The consistent need to balance model performance with privacy guarantees has led to research in the quantification of privacy and utility in LLMs. Sun and team (10.1145/3696410.3714531) focused on researching the effectiveness of privacy-preserving algorithms in LLMs, concluding that the CusText+ technique was optimal for robust privacy preservation. Das and team (das2025revisitingprivacyutilityefficiency) attempted to redefine privacy and utility specifically in the context of fine-tuning LLMs, evaluating different fine-tuning techniques such as differential privacy and LoRA. Pasch and team (pasch-cha-2025-balancing) studied anonymizations and utility in LLMs, concluding that anonymizations can reduce privacy leakage by 97-99% while only causing a 10% degradation in the model. These results serve as direct motivation for our work, as we choose to mask PIIs anonymously to reduce PII leakage while minimizing the impact on LLM performance.
Although existing studies have explored the impact of privacy-preserving techniques on LLM utility, they typically evaluate privacy and utility considerations separately rather than within an integrated analytical framework. To the best of our knowledge, this work represents the first systematic investigation of the privacy-utility tradeoff for LLM mitigation techniques under a unified framework that jointly characterizes both dimensions.
3 Dataset Used
We utilize a raw version of the Enron Email Dataset (shetty2004enron) to construct our training, validation, and testing splits. The dataset is divided into three partitions following a 100:10:1 ratio—the training set contains 10,000 datapoints, the validation set 1,000 datapoints, and the testing set 100 datapoints, all randomly sampled. Each datapoint corresponds to an individual email sample containing one or more personally identifiable email addresses.
To evaluate the two key aspects of our approach—memorization and performance—the experiment is structured into two evaluation phases:
  • Memorization Evaluation via Generation: We employ a set of 16,000 English prompts sampled from a publicly available web crawl. These prompts are designed to mimic realistic, open-ended user queries rather than explicitly targeted extraction prompts, ensuring the evaluation reflects general LLM usage patterns.
  • Perplexity Evaluation: To assess language modeling performance, we use the testing subset of the Enron dataset. This provides an accurate measure of how fine-tuned models generalize to unseen yet domain-consistent data.
For randomized masking, we construct a small, curated list of commonly used domain names and top-level domains (TLDs), publicly available on our GitHub repository  (joshi2024rmft_code). Each list contains approximately 15 examples, used to generate realistic but anonymized email variants during masking.
3.1 Exploratory Data Analysis
Before applying any dataset modifications, we conducted an exploratory data analysis (EDA) on the Enron Email dataset to better understand the statistical distribution of personally identifiable email addresses. Specifically, we aimed to quantify how often certain PIIs recur across the dataset, since our hypothesis—consistent with Kandpal and team  (kandpal-etal-2024-user) —is that frequent token repetition correlates with higher memorization likelihood in LLMs.
Figure 1 illustrates the distribution of the top ten most frequent email addresses across both the training and testing subsets. Some addresses appear nearly 900 times among the 10,000 datapoints in the training split, indicating strong duplication patterns. Such high-frequency occurrences are particularly problematic for privacy, as repeated exposure during fine-tuning can lead models to memorize entire token sequences corresponding to these PIIs. Additionally, we observed that several email addresses appear in both training and testing subsets, highlighting potential data leakage across splits.
These findings directly motivated our two dataset-level interventions: (1) Randomized Masked Fine-Tuning (RMFT), which reduces recurrence through randomized replacements while preserving structure; and (2) Deduplication, which explicitly removes repeated headers to prevent repeated exposure to the same PII tokens.
(a) Training set
(b) Testing set
Figure 1: Distribution of the top ten most frequent email addresses in training and testing sets.
4 Methodology
We perform experiments to test the effectiveness of randomized masking in preventing PII memorization. As we have taken our inspiration from Kandpal and team’s results on N-Gram deduplication, we also compare our results with an adapted version of deduplication for PIIs. Figure 2 depicts a summarized version of methodologies used in randomized masked fine-tuning and deduplication.
Figure 2: Methodology used for Randomized Masked Finetuning vs Deduplication. This diagram depicts how the index tables are formulated for either technique and how the updated index tables differ. If we have $N$ datapoints, $k$ emails occurring on each datapoint on average, and $u$ unique emails, then the number of times each email appears in the original index table is $Nk/u$. In RMFT, the updated index table maintains NkNk rows, but each email now appears only once and the number of unique emails increases. In deduplication, the updated index table has just $u$ rows with each email appearing once.
Our experiment compares randomized masked fine-tuning with deduplication. We have created a suite of evaluations that considers both memorization and performance of trained models. We use metrics defined in Section 4.2 to compare memorization and performance of the fine-tuning techniques.
Over both phases we are fine-tuning our model for 3 epochs, checkpointing the LLM at every 10% training steps, utilizing a fixed learning rate of 0.00002 over all three training methods, as inspired by Borkar’s work  (borkar2025privacyrippleeffectsadding). We also save the data seen until every checkpoint and compare checkpoint-based metric calculation for the different fine-tuning techniques. For generation experiments, greedy decoding with no sampling is used, setting the temperature to 0, top_p to 1.0, top_k to 0, and deterministic beam search switched off. The average TER, average SER, and average mean delta perplexity are calculated as the average respective values over the 30 checkpoints.
4.1 Deduplication vs RMFT
For the experiment comparing deduplication and RMFT, we evaluate on three criteria: memorization, model performance, and training performance. For training performance, time complexity is considered.
4.1.1 Randomized Masking
As defined in the introduction, RMFT masks multiple recurring PIIs with a randomized, structured PII that closely mimics the original PII. Using this methodology, we reduce the occurrences of certain PIIs from thousands down to a single occurrence in the dataset.
This is done using an indexed table of email addresses and their occurrences over the datapoints, their starting index and ending index. Each of these occurrences is sorted based on a combination of datapoint index and starting index. The first occurrence is left untouched, but the rest of the occurrences are masked with a randomized PII. For this experiment we utilize email addresses as the PII. Our randomized creation of emails is as follows:
  1. 1.
    We create a set of first names, last names, domain names, and TLDs as they appear in the set of originally available emails.
  2. 2.
    For each email to be masked, an anchor is decided randomly. The anchor is a part of the email address that remains unchanged.
  3. 3.
    We select the remaining two parts of the email address randomly from the collected sets of parts in Step 1.
  4. 4.
    The email address is created with the selected parts and anchor. The newly generated email replaces the original email at the exact position where it occurred.
For example, if the original email is kay.mann@enron.com, we might randomly select the last name as the anchor and replace the remaining components with sampled parts from the collected sets. This could yield bob.mann@yahoo.org, preserving structural consistency (FIRST.LAST@DOMAIN.TLD) while preventing exact memorization.
4.1.2 Deduplication Training
We adapt deduplication for PII-aware context, closely following the structure defined by Lee and team  (lee2022deduplicating). Lee and team define deduplication as dropping datapoints that contain token sequences of 400 tokens or more that are repeated multiple times. We adapt it to PII-aware context, where we drop datapoints that have a repeating email address.
Instead of discarding entire datapoints, we remove only the headers containing repeated email addresses, retaining the main email text. This preserves natural language context while still eliminating recurring PIIs from the structural header format. Such selective deduplication provides a more balanced tradeoff between privacy and linguistic richness.
4.2 Metrics Used
We use two metrics to calculate and compare memorization: Total Extraction Rate (TER) and Seen Extraction Rate (SER). We utilize Mean Delta Perplexity (MDP) to compare the performance of the training techniques.
4.2.1 Memorization Metrics
Total Extraction Rate is defined as the number of email addresses leaked at a checkpoint divided by the total number of email addresses available in the original set. Leaked email addresses are extracted from generated text using regex pattern matching for the format [first_name].[last_name]@[domain].[tld], with matches validated against the original email set.
TER is calculated as:
$TER_{ckpt_{i}}=\left(\frac{|LeakSet_{ckpt_{i}}|}{|OG\text{-}EMAIL\text{-}SET|}\right)\times 100$ (1)
Seen Extraction Rate is defined as the number of email addresses leaked at a checkpoint divided by the total number of emails seen until that checkpoint. To calculate the total number of emails seen until checkpoint $i$, we take the union of emails seen at checkpoint $i$ with the set of emails seen before checkpoint $i$.
Seen Extraction Rate is calculated as:
$SER_{ckpt_{i}}=\left(\frac{|LeakSet_{ckpt_{i}}|}{|SeenSet_{ckpt_{i}}\cup SeenSet_{ckpt_{i-1}}|}\right)\times 100$ (2)
Both TER and SER play an important role in determining memorization within the checkpoints of a given training technique. Higher SER and TER values mean a higher rate of emails are being memorized. Lower levels of TER and SER are preferred for better privacy.
4.2.2 Performance Metrics
We utilize perplexity on a fixed set of prompts for all checkpoints of the three training techniques to calculate the performance of each. We use the WikiText2 benchmark testing set, which contains around 2,500 prompts.
As the variance in calculated perplexities can be high, we utilize Mean Delta Perplexity (MDP), which measures the increase in perplexity compared to baseline.
Mean Delta Perplexity is calculated as:
$\Delta PP_{ckpt_{k}}=\sum_{i=1}^{N}\left((PP_{ckpt_{k}}^{TM})_{p_{i}}-(PP_{ckpt_{k}}^{BL})_{p_{i}}\right)\times\frac{1}{N}$ (3)
This calculates the average difference in perplexity per prompt for a given checkpoint in two different training methods. For this experiment, we associate a higher MDP with worse performance, considering baseline perplexity as our standard threshold.
5 Results and Observations
In this section, we summarize the key quantitative outcomes of our experiments across three fine-tuning techniques: baseline fine-tuning, randomized masked fine-tuning (RMFT), and deduplication training. Metrics are averaged over 30 checkpoints unless otherwise stated. Table 1 summarizes the memorization and perplexity observations.
Table 1: Summary of observations over three different training techniques. Each value represents an average over all checkpoints. Extraction rates are percentages. The best performing technique in each column is bolded.
Training PPL TER SER
Technique ($\downarrow$) ($\downarrow$) ($\downarrow$)
Baseline 6.432 0.308 0.342
RMFT 6.798 0.049 0.059
Deduplication 7.963 0.084 0.102
5.1 Extraction Rates
Both RMFT and deduplication significantly reduce memorization relative to baseline fine-tuning. RMFT achieves an 80.81% reduction in Total Extraction Rate (TER) and an 80.17% reduction in Seen Extraction Rate (SER) compared to the baseline, outperforming deduplication’s 69.38% TER and 68.84% SER reductions. This demonstrates RMFT’s superior ability to mitigate PII memorization while maintaining dataset utility.
At the final checkpoint, RMFT achieves an 87.62% reduction in TER and 86.0% in SER, compared to deduplication’s 74.22% and 72.9%, respectively. These results suggest that RMFT continues to improve privacy robustness over training progression. Figure 4 shows the Total Extraction Rates and Seen Extraction Rates for all three training techniques on a fixed web crawl dataset of 16,000 prompts.
(a) Total Extraction Rate (TER)
(b) Seen Extraction Rate (SER)
Figure 3: Total Extraction Rate (TER) and Seen Extraction Rate (SER) across training checkpoints for all three fine-tuning techniques. Lower values indicate fewer PIIs memorized by the model.
5.2 Mean Delta Perplexity
The Mean Delta Perplexity (MDP) curves exhibit a consistent pattern: RMFT has consistently less degradation in performance compared to deduplication. RMFT achieves a significantly lower average perplexity increase (+5.73%) relative to the baseline, compared to deduplication’s +23.85%. RMFT maintains 14.5% lower average perplexity on all checkpoints than deduplication, confirming reduced performance degradation. The final training checkpoint for RMFT has a 5.56% increase in perplexity compared to 26% increase in deduplication. Figure 4(a) shows the Mean Delta Perplexity plot for both RMFT and deduplication at each checkpoint.
(a) Mean Delta Perplexity
(b) Average Perplexity per Checkpoint
Figure 4: (a) Mean Delta Perplexity of Deduplication vs Randomized Masked Finetuning (RMFT). Lower MDP indicates less deviation from baseline performance. (b) Average Perplexity per Checkpoint for all training techniques. This curve conforms the Mean Delta Perplexity of RMFT being lower than Deduplication
5.3 Time Complexity Analysis
We calculate the time complexity of training and compare the two techniques—deduplication and RMFT. We only consider the time complexity of preprocessing, as both training techniques follow the same training steps once the data is preprocessed.
We assume $N$ datapoints for training, each datapoint has on average $k$ PIIs appearing, the length of each datapoint is $l$, and the number of unique emails is $u$. The total number of PIIs occurring in the whole dataset would be $N\cdot k$.
5.3.1 Randomized Masked Finetuning
The time complexity for randomized masked fine-tuning is $O(N\cdot k\cdot|\log u|)$. The $|\log u|$ term arises because we sort the PIIs based on their appearance. This is important as we want to keep the first appearance untouched but need to mask the rest.
5.3.2 Deduplication
Our modified deduplication approach is divided into two phases: deduplication of headers and recombination of headers. The deduplication of headers takes $O(Nk)$ time. The recombination process takes $O(N\log N)$ time, and hence we report the dominating term as the final time complexity: $O(N\log N)$.
For our study, we took 10,000 training datapoints, each having 6 emails on average and 24,480 unique email addresses. RMFT takes 15 minutes and 25 seconds of preprocessing on average, whereas our modified deduplication takes 10 minutes and 12 seconds of preprocessing on an NVIDIA H200 GPU. This clearly indicates that deduplication is more efficient than randomized masked fine-tuning for preprocessing time.
6 Analysis
We perform a Pareto frontier analysis on our experimental results to identify the most optimal training checkpoint for a given percentage threshold increase in Mean Delta Perplexity (MDP) that yields the maximum percentage reduction in Total Extraction Rate (TER). We adopt TER percentage drop rather than SER percentage drop for this analysis, since the Pareto frontier captures a global optimal tradeoff between privacy and performance. TER reflects a global measure of overall memorization across the dataset, whereas SER reflects localized memorization behavior limited to the model’s seen subset.
We define a MaxTER function that returns the maximum TER percentage drop for a given threshold of MDP:
$MaxTER_{t}(\tau)=\max_{\{i\mid MDP_{i,t}\leq\tau\}}\Delta TER_{i,t}$ (4)
where $\tau$ is the MDP threshold acceptable, $i$ is the checkpoint number, and $t$ is the training technique.
Percentage MDP is calculated as:
$MDP_{i,t}\%=\left(\frac{PPL_{i,t}-PPL_{i,b}}{PPL_{i,b}}\right)\times 100$ (5)
Percentage TER Drop is calculated as:
$\Delta TER_{i,t}\%=\frac{TER_{i,t}-TER_{i,b}}{TER_{i,b}}\times 100$ (6)
Figure 5 presents the MaxTER analysis comparing deduplication and RMFT. Notably, deduplication exhibits no feasible checkpoints below a 20% increase in MDP, marking the region prior to this threshold as infeasible. In contrast, RMFT demonstrates consistent performance across all MDP thresholds, achieving higher reductions in TER for equivalent or lower increases in perplexity. We compute the Area Under the Response Curve (AURC) to quantify overall privacy-performance efficiency. RMFT achieves an AURC of approximately 92.52 units², substantially surpassing deduplication’s 17.09 units².
Figure 5: MaxTER analysis identifying checkpoints that optimize the tradeoff between TER reduction and perplexity increase. RMFT outperforms deduplication by achieving greater TER reductions while incurring lower perplexity penalties. No deduplication checkpoint satisfies the constraint of perplexity increase below 20% (threshold $\tau$), representing the grey infeasible region
This analysis addresses how to evaluate the tradeoff between reduced memorization and potential degradation in model performance. The proposed MaxTER function provides an effective framework for identifying the optimal checkpoint for each fine-tuning technique by quantifying the balance between reduced memorization (percentage drop in TER) and increased model perplexity (percentage rise in MDP). RMFT consistently outperforms deduplication in memorization reduction, achieving lower TER (0.049 vs 0.084) and SER (0.059 vs 0.102) across all evaluated conditions. On in-distribution evaluation using the Enron held-out test set, RMFT achieves substantially superior AURC (92.52 vs 17.09 units²), demonstrating its effectiveness when deployment conditions match training distributions—the primary use case for privacy-preserving fine-tuning.
7 Conclusion
In conclusion, this study demonstrates that Randomized Masked Fine-tuning (RMFT) is an effective and computationally efficient approach to mitigating the memorization of personally identifiable information (PIIs) in large language models. Compared to traditional deduplication, RMFT achieves substantially greater reductions in Total and Seen Extraction Rates while maintaining lower degradation in model performance as measured by Mean Delta Perplexity. Through MaxTER analysis, we show that RMFT consistently provides a more favorable tradeoff between privacy preservation and performance retention. Our results establish RMFT as a promising, scalable fine-tuning strategy for enhancing privacy robustness in LLMs when deployment distributions match training conditions without compromising their generative quality—ideal for privacy-preserving training.
8 Limitations and Future Work
While RMFT demonstrates strong potential as a privacy-preserving fine-tuning strategy, several limitations remain. First, our experiments are restricted to the fine-tuning stage of model training, whereas the RMFT framework can, in principle, be applied to pretraining as well. The effect of randomized masking during large-scale pretraining—where token frequency and contextual learning dynamics differ significantly—remains unexplored and may yield different outcomes in terms of both privacy and performance.
Finally, RMFT’s current success is closely tied to the structured nature of PIIs such as email addresses, which can be effectively masked with realistic, randomized replacements. Extending this method to mitigate memorization of arbitrary token sequences—where semantic and syntactic coherence must be preserved—poses a greater challenge. This limitation also motivated our adaptation of deduplication into a PII-aware variant, as standard deduplication better addresses generic token-level repetition.
Acknowledgements
We thank the reviewers for their valuable feedback.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning with respect to privacy protection in large language models. There are many potential societal consequences of our work, particularly in enabling safer deployment of LLMs in privacy-sensitive domains such as healthcare, finance, and legal services. By reducing PII memorization, our approach can help mitigate privacy risks while maintaining model utility, contributing to more responsible AI development practices.
Appendix A Example of Randomized PII Creation
Figure 6 shows how the randomizer in RMFT creates new email addresses. The index table at the top is created using the original dataset. The first name store and last name store are populated using the already existing emails in the dataset. The domain store contains popularly used domains.
Figure 6: This diagram shows an example of how RMFT replaces an occurrence of an email with a randomized email. The randomizer selects an anchor (first name, last name, or domain) and keeps it fixed while selecting the remaining parts randomly from the store.
A.1 Example Dataset
Consider the following email:
Message-ID: <33099599.1075845686859.JavaMail.evans@thyme>
Date: Wed, 4 Apr 2001 02:32:00 -0700 (PDT)
From: kay.mann@enron.com
To: suzanne.adams@att.net
Subject: Re: Wednesday
How’s everything coming up? Did Warren help you last night?
Kay
Let’s take the example of suzanne.adams@att.net:
  1. 1.
    The three facets of the email are: suzanne (first name), adams (last name), and att.net (domain).
  2. 2.
    The randomizer selects an anchor randomly, say first name.
  3. 3.
    The randomizer selects random parts for last name and domain, say watterberg and google.com.
  4. 4.
    The new email would be suzanne.watterberg@google.com.
Appendix B MaxTER Evaluation on Different Datasets
We present evaluations of RMFT versus deduplication via the MaxTER function on three datasets: Enron held-out test dataset, WikiText2 test dataset  (radford2019language), and randomly collected web crawl prompt dataset. Table 2 summarizes all results.
(a) Enron Held-Out Dataset
(b) WikiText2 Test Dataset
(c) Web Crawl Prompts
Figure 7: MaxTER curves for RMFT vs Deduplication across three different datasets. The curves show the privacy-performance tradeoff measured by Area Under the Response Curve (AURC).
Table 2: Comparison of training techniques across three different datasets. MaxTER AURC represents the direct relation between memorization reduction and performance.
Enron Held Out WikiText2 Web Crawl
Technique Avg PPL AURC Avg PPL AURC Avg PPL AURC
Baseline 6.432 - 605.23 - 2168.39 -
RMFT 6.798 92.52 585.43 75.548 2214.70 76.84
Dedup 7.963 17.09 531.97 82.081 1847.56 82.09
From the above table we can see how deduplication has a higher MaxTER AURC for different datasets such as WikiText2 test and web crawl prompts, whereas RMFT maintains a higher MaxTER curve for Enron held-out dataset. RMFT heavily outperforms deduplication in the Enron held-out dataset, whereas the performance on WikiText2 is comparable between both techniques. Deduplication outperforms RMFT on randomly sampled prompts from web crawl.
Appendix C Time Complexity Analysis
For this analysis, we assume that we have $N$ datapoints, each datapoint being $l$ length long, each datapoint having $k$ emails on average, and there are $u$ unique emails all over the dataset.
C.1 Randomized Masked Finetuning
We need to sort all email addresses fetched in the index table based on their occurrence. Each email appears $\frac{N\cdot k}{u}$ times. Therefore, using quick sort the time complexity for sorting becomes:
$TC_{sorting}$ $=\frac{N\cdot k}{u}\log\left(\frac{N\cdot k}{u}\right)$ (7)
$=\frac{N\cdot k}{u}(\log N+\log k-\log u)$ (8)
As we repeat this for $u$ unique emails:
$TC_{sorting}=u\cdot\frac{N\cdot k}{u}(\log N+\log k-\log u)$ (9)
In our case the number of unique emails is of higher magnitude than number of datapoints $N$ and average email appearance $k$. Therefore:
$TC_{sorting}=O(N\cdot k\cdot|\log u|)$ (10)
The randomization and replacement of PIIs occur in $O(1)$ for $N\cdot k$ emails:
$TC_{RR}=O(N\cdot k)$ (11)
The time complexity of RMFT is:
$\mathbf{TC_{RMFT}=O(N\cdot k\cdot|\log u|)}$ (12)
C.2 Deduplication
The deduplicator takes $O(1)$ time for deduplication and we carry it out for $u$ unique emails:
$TC_{Deduplicator}=O(u)$ (13)
We have $N$ headers on average and sorting them gives:
$TC_{HeaderSorting}=O(N\log N)$ (14)
The remerging takes $O(N)$. Therefore:
$TC_{Deduplication}=O(u)+O(N\log N)+O(N)$ (15)
We report the dominating term:
$\mathbf{TC_{Deduplication}=O(N\log N)}$ (16)